OpenStudy (anonymous):

A box is has dimensions of 14 inches long, 1.5 feet wide, and 7 inches high. What is the volume of the box? The formula for the volume is V = l • w • h

OpenStudy (anonymous):

it is telling you to multiply right?

OpenStudy (anonymous):

so your job is to compute \[14\times 1.5\times 12\times 7\]

OpenStudy (anonymous):

the 12 is in there because there are 12 inches in a foot, so the number of inches in 1.5 feet is \(1.5\times 12\)